Suppose ACT Reading scores are normally distributed with a mean of 21.4 and a standard deviation of 6.4. A university plans to award scholarships to students whose scores are in the top 5%. What is the minimum score required for the scholarship? Round your answer to the nearest tenth, if necessary
Answer by Boreal(15235)

The z-score for 0.95 is +1.645
z=(x-mean)/sd
1.645*6.4=x-21.4
x=21.4+(6.4*1.645)=21.4+10.5=31.9
The 95th percentile is at 31.9, and that is the minimum score.